# A Unified Approach to Routing and Cascading for LLMs

### v1

@@ 4 Cascade Routing as Cascade Generalization
Both routing and cascading are powerful techniques that enable the efficient use of multiple models. However, their use is often orthogonal: while routing is useful when we have several specialized models that are experts at specific tasks, cascading is more beneficial when input queries vary in difficulty. In this section, we therefore present cascade routing, which is a generalization of both techniques. Proofs for all theorems and lemmas in this section are included in App. C.
@@
While cascade routing extends cascading and can therefore use the same hyperparameter optimization scheme, it also introduces additional challenges which we address in the following paragraphs.
@@
5 Evaluation
We evaluate the performance of cascade routing and demonstrate that it significantly outperforms all other strategies. Additionally, we show that our new cascading approach surpasses the threshold-based cascading method (see Corollary 1). For this purpose, we first conduct experiments on RouterBench (Hu et al., 2024), a benchmark specifically designed to evaluate routing and cascading (Section 5.1). Next, we test cascade routing on two additional benchmarks to evaluate its performance in more realistic scenarios (Section 5.2). Lastly, we perform an ablation study to examine the impact of various design choices in cascade routing on performance and runtime (Section 5.3).
5.1 RouterBench
RouterBench (Hu et al., 2024) is a benchmark developed to evaluate the efficacy of different model selection strategies. It includes questions from seven diverse benchmarks, such as MMLU (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), and MBPP (Austin et al., 2021), alongside answers from eleven different models ranging from GPT-4 (OpenAI, 2023) to Mistral-7B (Jiang et al., 2023). All models are evaluated under both zero-shot and five-shot settings.
Quality and Cost Estimates
Similar to (Hu et al., 2024), we estimate quality and cost by adding zero-centered Gaussian noise to their true values. Both cost and quality estimates are modeled as linear functions fitted on these noisy signals. We define the variance of the noisy signal as $\sigma_{\text{before}}^{2}$ before model computation and $\sigma_{\text{after}}^{2}$ after. By ensuring that $\sigma_{\text{after}}<\sigma_{\text{before}}$, this setup reflects the increased accuracy in cost and quality estimates after model computation, which is an essential requirement for cascading to perform well. To explore different uncertainty levels, we vary the variances to simulate low-, medium-, and high-noise scenarios, with exact values for the variances detailed in App. D.
Models
We evaluate cascade routing on RouterBench using three, five, and all eleven models available for model selection, ensuring a comprehensive evaluation across a range of scenarios. The exact models for each scenario are provided in App. D.
Figure 2: Quality-cost tradeoff on RouterBench for five models and medium-noise.
Strategies
We compare cascade routing against several baseline strategies, including the routing strategy described in Section 2, the threshold-based cascading approach from prior work (Corollary 1), and the optimal cascading strategy (Theorem 2). Additionally, as in (Hu et al., 2024), we include a baseline that linearly interpolates cost and quality on the Pareto frontier of the models. Fig. 2 illustrates the performance of these strategies with five models in a medium-noise setting.
Evaluation Metric
For each method, we evaluate performance using cost budgets ranging from the cheapest to the most expensive model, as shown in Fig. 2. This produces a quality-cost curve for each strategy. Following (Hu et al., 2024), we use the Area Under the Curve (AUC) as the performance metric.
@@ 5.1 RouterBench
Cascade (Ours) 81.129 17281.12917281.129\,17281.129 172 76.100 59176.10059176.100\,59176.100 591 72.666 79172.66679172.666\,79172.666 791 83.049 49883.04949883.049\,49883.049 498 75.148 59975.14859975.148\,59975.148 599 70.175 50570.17550570.175\,50570.175 505 84.449 86384.44986384.449\,86384.449 863 75.103 98275.10398275.103\,98275.103 982 70.255 39270.25539270.255\,39270.255 392
Cascade Routing 82.339 79482.33979482.339\,79482.339 794 76.556 64476.55664476.556\,64476.556 644 73.232 31873.23231873.232\,31873.232 318 84.342 56284.34256284.342\,56284.342 562 76.315 41976.31541976.315\,41976.315 419 72.744 80172.74480172.744\,80172.744 801 87.275 86387.27586387.275\,86387.275 863 77.615 53377.61553377.615\,53377.615 533 74.397 70974.39770974.397\,70974.397 709
[…]
@@ Results
Table 1 presents the results for the zero-shot setting, with the five-shot results detailed in App. E. Cascade routing consistently outperforms all baseline strategies with performance gains between $1\%$ to $4\%$, which measured relatively to the naive linear interpolation baseline means that cascade routing improves by $13\%$ to $80\%$ over the baselines. This performance gap widens as more models are available and narrows under higher noise levels, indicating that cascade routing is most effective with large model sets and accurate cost and quality estimates. Furthermore, our new cascading strategy outperforms the threshold-based cascade by up to $2\%$, reinforcing the practical relevance of our theoretical results.
5.2 Other Benchmarks
RouterBench does not provide log probabilities associated with model answers, which constrains the construction of more realistic quality estimates using features like perplexity. To address this, we develop two benchmarks that better simulate practical use cases for cascade routing.
Datasets
We perform experiments on classification and open-form reasoning tasks. The classification benchmarks include ARC-Challenge (Clark et al., 2018), MMLU-Pro (Wang et al., 2024), and MixEval (Ni et al., 2024). For open-form reasoning tasks, we use MMLU-Pro and GSM8k (Cobbe et al., 2021). In classification, models select a single option representing their answer, with no intermediate reasoning process. In contrast, open-form reasoning allows models to generate their answers after reasoning. We split each benchmark into a training set for training quality and cost estimates and a test set for evaluation. We evaluate nine models, three each from the Llama-3.1 (AI@Meta, 2024), Gemma (Gemma Team et al., 2024), and Mistral (Jiang et al., 2023) model families. The exact models are specified in App. D.
Quality and Cost Estimates
We estimate model quality by fitting a linear model based on features that reflect model uncertainty, such as perplexity (Gupta et al., 2024). Further features include the originating benchmark of each sample and whether earlier models in the cascade agree on their prediction. Full details on the features are provided in App. D. The linear model is trained using these features on the benchmark’s training split.
Classification Open-Form
Llama Gemma Mistral Llama Gemma Mistral
Linear Interp. 74.282 60074.28260074.282\,60074.282 600 61.683 00761.68300761.683\,00761.683 007 63.389 43563.38943563.389\,43563.389 435 79.113 00079.11300079.113\,00079.113 000 54.097 63654.09763654.097\,63654.097 636 53.859 56753.85956753.859\,56753.859 567
Routing 74.922 47474.92247474.922\,47474.922 474 64.439 96464.43996464.439\,96464.439 964 64.890 30064.89030064.890\,30064.890 300 79.318 91079.31891079.318\,91079.318 910 58.395 59158.39559158.395\,59158.395 591 58.705 10758.70510758.705\,10758.705 107
Cascade (Baseline) 74.791 91774.79191774.791\,91774.791 917 54.308 09354.30809354.308\,09354.308 093 61.216 45061.21645061.216\,45061.216 450 79.232 86279.23286279.232\,86279.232 862 56.175 06556.17506556.175\,06556.175 065 48.288 64648.28864648.288\,64648.288 646
Cascade (Ours) 75.424 17275.42417275.424\,17275.424 172 62.817 29962.81729962.817\,29962.817 299 63.360 21763.36021763.360\,21763.360 217 79.683 35579.68335579.683\,35579.683 355 57.672 17757.67217757.672\,17757.672 177 55.507 38055.50738055.507\,38055.507 380
Cascade Routing 75.523 23075.52323075.523\,23075.523 230 64.699 68564.69968564.699\,68564.699 685 64.970 63464.97063464.970\,63464.970 634 79.837 04179.83704179.837\,04179.837 041 59.615 95659.61595659.615\,95659.615 956 58.689 94958.68994958.689\,94958.689 949
Table 2: AUC scores on a classification and open-form reasoning benchmark.
@@
Table 2 presents the results for the Llama, Gemma, and Mistral model families across both benchmarks. Cascade routing consistently performs on par with or outperforms all baselines, though with narrower margins reaching up to $1.2\%$. This reduced gain can be attributed to two main factors. First, the quality and cost estimates are very noisy, leading to performance gains over the naive baseline similar to those observed in high-noise scenarios on RouterBench. Second, the cascading strategy sometimes underperforms compared to the linear interpolation baseline, indicating that the post-computation features used for quality estimation, such as perplexity, do not provide sufficient advantage to warrant running models in a cascading fashion.
As expected, cascade routing is most effective when both routing and cascading outperform the linear interpolation baseline. When cascading offers no or very small performance improvement, cascade routing typically reduces to pure routing. Larger performance gains are observed only when cascading adds value beyond routing. Our results suggest that more sophisticated methods are needed to enhance quality estimates and improve the overall effectiveness of all model selection strategies in practical applications.
5.3 Ablation Study
We conduct an ablation study to examine the impact of various design choices in cascade routing on performance and runtime. Runtime is a critical factor because the overhead introduced by the strategy must be negligible compared to the time required for model computation. If the strategy adds significant overhead, its performance gains may be offset by the increased runtime.
Low-Noise Medium-Noise High-Noise
AUC (%) Time (ms) AUC (%) Time (ms) AUC (%) Time (ms)
Cascade Routing 87.262 69887.26269887.262\,69887.262 698 12.047 85812.04785812.047\,85812.047 858 77.604 61977.60461977.604\,61977.604 619 12.754 49712.75449712.754\,49712.754 497 74.407 47274.40747274.407\,47274.407 472 9.412 8919.4128919.412\,8919.412 891
Slow 87.293 51187.29351187.293\,51187.293 511 80.037 32280.03732280.037\,32280.037 322 77.624 59177.62459177.624\,59177.624 591 87.231 79987.23179987.231\,79987.231 799 74.407 20474.40720474.407\,20474.407 204 67.450 26567.45026567.450\,26567.450 265
Greedy 85.932 29185.93229185.932\,29185.932 291 1.635 9611.6359611.635\,9611.635 961 77.162 71177.16271177.162\,71177.162 711 1.587 7611.5877611.587\,7611.587 761 74.351 52874.35152874.351\,52874.351 528 1.040 7491.0407491.040\,7491.040 749
No-Expect 85.977 67585.97767585.977\,67585.977 675 3.370 $3$.3703333.370\,$3$.370 333 77.111 51377.11151377.111\,51377.111 513 3.039 9083.0399083.039\,9083.039 908 74.347 16374.34716374.347\,16374.347 163 1.773 9991.7739991.773\,9991.773 999
Table 3: AUC scores and average runtime for various variations of cascade routing on RouterBench when using all eleven models.
@@
Table 3 presents the results of the ablation study. As expected, the Slow variation is almost an order of magnitude slower while achieving similar performance. In contrast, both Greedy and No-Expect are faster but perform worse in the low- and medium-noise scenarios by $0.5\%$ to $1.3\%$. Interestingly, there seems to be a much smaller performance gap in the high-noise scenario. This is due to the very low variance in the quality estimates, since the linear model used for quality estimation predicts an almost constant value for each query in this scenario, making the expected value computation less important.
@@ 6 Related Work
Routing is a widely studied problem in machine learning, particularly in the task of directing input queries to specialized expert models. One of the most common applications of routing is model selection for natural language input queries with a known correct answer (Ding et al., 2024; Hari and Thomson, 2023; Liu et al., 2024; Jang et al., 2023; Nguyen et al., 2024; Sakota et al., 2024; Shnitzer et al., 2023). All these works train a machine learning model to predict whether a given model will correctly answer a query. Though the setups in these works are largely similar, they vary in certain specifics, such as the type of input queries or the features used for quality estimation.
Routing is also applied in other areas. For instance, Lu et al. (2024); Ong et al. (2024) use preference data to train a quality estimator, which facilitates routing in scenarios involving real-world user queries where clear ground-truth answers may not exist. Additionally, Chen et al. (2022) employ routing for API selection in multi-label classification tasks, focusing on directing queries to the appropriate API based on task requirements. Similarly, Zhang et al. (2024) apply routing in software agent environments, directing user issues to the agent most suited to handle them.
@@
Cascading techniques are primarily used to reduce inference costs by employing smaller models initially and only cascading to larger models if the smaller ones fail to provide a sufficiently accurate answer. Most often, cascading decisions are based on the smaller model’s confidence in its own predictions (Chen et al., 2023; Ramírez et al., 2024; Varshney and Baral, 2022). However, alternative techniques also exist. For example, Madaan et al. (2023) propose running models multiple times and measuring the variance in their responses to decide whether to cascade to a larger model.
@@
There has also been specific research on quality estimation within cascading frameworks. Gupta et al. (2024) examine various measures of uncertainty in language model answers, evaluating their impact on cascading performance. Meanwhile, Jitkrittum et al. (2023) explore failure cases in cascading mechanisms that rely on uncertainty, introducing alternative quality measures that enhance cascade efficiency. Lastly, Xue et al. (2023) apply cascading to majority voting for a single model to obtain a method called dynamic voting: the cascade stops depending on the aggregated answers of all previous model computations. This allows the system to process simpler queries using fewer votes while allocating more computational resources for harder queries.
7 Limitations
While cascade routing provides a promising approach to improve model selection strategies, it requires accurate cost and quality estimates to be effective. However, in Section 5.2 we found that quality estimates based on current state-of-the-art methods contain significant noise. This limits the performance improvement of cascade routing over routing, as the quality estimates are not accurate enough to guide the selection of the best model. As shown in Section 5.1, lower noise in the quality and cost estimates lead to much better performance of all model selection strategies, indicating that further work that aims to reduce the noise in the quality and cost estimates is needed to fully leverage the potential of cascade routing.
Another limitation is the necessity of a training dataset to optimize the hyperparameters associated with cascade routing. While both routing and cascading suffer from the same disadvantage, the training data required for routing is much more minimal, since it only needs to estimate two hyperparameters. In contrast, cascade routing requires estimating multiple hyperparameters on a more complex optimization surface.

### v2

@@ 4 Cascade Routing as Cascade Generalization
Both routing and cascading are powerful techniques that enable the efficient use of multiple models. However, their use is often orthogonal: while routing is useful when ex-ante quality estimates are accurate, cascading is more beneficial when post-hoc estimates are accurate. We therefore present cascade routing, which is a generalization of both techniques. Proofs for all theorems and lemmas in this section are included in App. C.
Brief Overview
In this section, we first define cascade routing and explain how it generalizes both routing and cascading. We then derive the optimal cascade routing strategy and solve several of the additional challenges that arise when applying cascade routing in practice.
@@
While cascade routing extends cascading and can therefore use the same hyperparameter optimization scheme, it also introduces additional challenges which we address now.
@@
[…]
@@
Cascade (Ours) 81.087 63781.08763781.087\,63781.087 637 76.163 85076.16385076.163\,85076.163 850 72.668 42972.66842972.668\,42972.668 429 83.057 02683.05702683.057\,02683.057 026 75.169 22175.16922175.169\,22175.169 221 70.179 55670.17955670.179\,55670.179 556 84.468 95884.46895884.468\,95884.468 958 75.095 88075.09588075.095\,88075.095 880 70.255 07870.25507870.255\,07870.255 078
Cascade Routing (Ours) 82.364 41682.36441682.364\,41682.364 416 76.553 92276.55392276.553\,92276.553 922 73.216 01573.21601573.216\,01573.216 015 84.329 59884.32959884.329\,59884.329 598 76.313 68376.31368376.313\,68376.313 683 72.745 10372.74510372.745\,10372.745 103 87.243 45687.24345687.243\,45687.243 456 77.568 91777.56891777.568\,91777.568 917 74.404 52174.40452174.404\,52174.404 521
5 Experimental Evaluation
We now evaluate the performance of cascade routing and demonstrate that it significantly outperforms all other strategies. Additionally, we show that our new cascading approach outperforms the threshold-based cascading method. For this purpose, we first conduct experiments on RouterBench (Hu et al., 2024), a benchmark specifically designed to evaluate routing and cascading (Section 5.1). Next, we test cascade routing on several additional benchmarks to evaluate its performance in more realistic scenarios (Section 5.2). In the appendix, we additionally perform an ablation study to examine the impact of various design choices in cascade routing on performance and runtime (App. F).
5.1 RouterBench
RouterBench (Hu et al., 2024) is a benchmark developed to evaluate the efficacy of different model selection strategies. It includes questions from seven diverse benchmarks, such as MMLU (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), and MBPP (Austin et al., 2021), alongside answers from eleven different models ranging from GPT-4 (OpenAI, 2023) to Mistral-7B (Jiang et al., 2023).
Quality and Cost Estimates
Similar to (Hu et al., 2024), we estimate quality and cost by adding zero-centered Gaussian noise to their true values. Both cost and quality estimates are modeled as linear functions fitted on these noisy signals. Thus, the quality estimate can be expressed as $\hat{q}_{W,b}(x)=W(q(x)+\epsilon)+b$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. A similar expression holds for the cost estimate. We define the variance of the noisy signal as $\sigma_{\text{ante}}^{2}$ before model computation (ex-ante estimates) and $\sigma_{\text{post}}^{2}$ after (post-hoc estimates). To explore different uncertainty levels, we vary the variances to simulate low-, medium-, and high-noise scenarios, with exact values for the variances given in Section D.1.
Models
We evaluate cascade routing on RouterBench using three, five, and eleven models available for model selection, ensuring a comprehensive evaluation across a range of scenarios. The exact models are provided in Section D.1.
((a)) Comparison of cascade routing with routing.
((b)) Comparison of cascade routing with cascading.
Figure 2: Difference in AUC performance between cascade routing and baseline strategies on RouterBench for various noise values. Red indicates cascade routing is much better, while blue indicates it is only a bit better.
Strategies
We compare cascade routing against several baseline strategies, including the routing strategy described in Section 2, the threshold-based cascading approach from prior work (Corollary 1), and the optimal cascading strategy (Theorem 2). Additionally, as in (Hu et al., 2024), we include a baseline that linearly interpolates cost and quality on the Pareto frontier of the models.
Evaluation Metric
For each method, we evaluate performance using cost budgets ranging from the cheapest to the most expensive model. This produces a quality-cost curve for each strategy. Following (Hu et al., 2024), we use the Area Under the Curve (AUC) as the performance metric.
@@ Results
Table 1 presents the results for the zero-shot setting, with the five-shot results detailed in App. H. Cascade routing consistently outperforms all baseline strategies with performance gains between $1\%$ to $4\%$, which measured relatively to the naive linear interpolation baseline means that cascade routing improves by $13\%$ to $80\%$ over the baselines. This performance gap widens as more models are available and narrows under higher noise levels, indicating that cascade routing is most effective with large model sets and accurate cost and quality estimates. Furthermore, our new cascading strategy outperforms the threshold-based cascade by up to $2\%$, reinforcing the practical relevance of our theoretical results.
Quality Estimation
To better understand the impact of quality estimation on model selection strategies, we additionally conduct experiments with five models under a broader range of varying noise levels. Fig. 2 illustrates the difference in AUC performance between cascade routing and baseline strategies for all possible noise levels. The results demonstrate that cascade routing consistently outperforms the baselines, achieving up to an $8\%$ improvement for cascading and up to a $12\%$ improvement for routing. Notably, the performance gap highlights key differences between the cascading and routing strategies. For routing, the value of $\sigma_{\text{ante}}$ is critical—high $\sigma_{\text{ante}}$ significantly reduces performance compared to cascade routing. Conversely, for cascading, $\sigma_{\text{post}}$ plays a more influential role, with higher values causing substantial performance degradation. These findings underscore the importance of accurate quality estimation for both strategies. Cascade routing proves to be a more robust solution by unifying the strengths of both approaches and effectively leveraging low $\sigma_{\text{ante}}$ and low $\sigma_{\text{post}}$ to enhance performance.
5.2 Real-World Benchmarks
We now show that cascade routing outperforms baselines on more realistic benchmarks with quality estimates that can be used in real-world scenarios. We differentiate in our analysis between benchmarks where accurate quality estimation is available and those where it is not.
Accurate Quality Estimation
We first evaluate cascade routing on two benchmarks that allow for accurate quality estimation. First, in the domain of software engineering, it is often easier to generate tests to reproduce specific issues than to fix them. We therefore use SWE-Bench (Jimenez et al., 2024) as a benchmark where accurate post-hoc quality estimation is available. Specifically, we assume that the quality of a model’s response can be accurately estimated by testing it on the ground-truth test cases. Second, to simulate a use-case where ex-ante quality estimation is accurate, we use the Math and Coder models from the Qwen-2.5 model family (Yang et al., 2024; Hui et al., 2024) and evaluate them on a combination of Minerva Math (Lewkowycz et al., 2022) and LiveCodeBench (Jain et al., 2024). To obtain accurate quality estimates, we incorporate a sample’s origin benchmark as a feature in the quality estimation model. For all details about the benchmarks, models, and estimators, we refer to Section D.1.
Table 2: AUC scores on practical benchmarks. On the left, resp. right, side we show the benchmarks with good, resp. poor, quality estimates. The highest numbers are bolded, and underlined numbers are within the $95\%$ confidence intervals of the highest number. For a discussion on confidence intervals, refer to App. E. In App. G, we present benchmark-specific AUC values for results averaged over several benchmarks.
SWE-Bench Math+Code Classification Open-Form
10 Models 5 Models Qwen Llama Gemma Llama Gemma
Linear Interp. 40.505 39240.50539240.505\,39240.505 392 38.637 49438.63749438.637\,49438.637 494 39.626 90139.62690139.626\,90139.626 901 74.282 60074.28260074.282\,60074.282 600 61.683 00761.68300761.683\,00761.683 007 79.113 00079.11300079.113\,00079.113 000 54.097 63654.09763654.097\,63654.097 636
Routing 40.470 85440.47085440.470\,85440.470 854 39.395 79639.39579639.395\,79639.395 796 47.463 30747.46330747.463\,30747.463 307 74.922 47474.92247474.922\,47474.922 474 $64.44$ 79.318 91079.31891079.318\,91079.318 910 58.395 59158.39559158.395\,59158.395 591
Cascade (Baseline) 38.519 80238.51980238.519\,80238.519 802 45.890 31545.89031545.890\,31545.890 315 37.677 35537.67735537.677\,35537.677 355 74.807 58274.80758274.807\,58274.807 582 54.319 90154.31990154.319\,90154.319 901 79.232 86279.23286279.232\,86279.232 862 56.175 06556.17506556.175\,06556.175 065
Cascade (Ours) $53.20$ $50.94$ 46.763 93446.76393446.763\,93446.763 934 $75.46$ 62.786 11862.78611862.786\,11862.786 118 79.224 97879.22497879.224\,97879.224 978 56.184 75556.18475556.184\,75556.184 755
Cascade Routing (Ours) 54.121 85254.12185254.121\,85254.121 852 51.090 58651.09058651.090\,58651.090 586 48.549 85448.54985448.549\,85448.549 854 75.516 72875.51672875.516\,72875.516 728 64.835 86164.83586164.835\,86164.835 861 79.877 36379.87736379.877\,36379.877 363 59.656 50359.65650359.656\,50359.656 503
@@
Table 2 (left) shows the results for both benchmarks. In SWE-Bench, our methods outperform baseline strategies by up to $14\%$. As expected, the routing strategy does not outperform the trivial baseline on this benchmark, as ex-ante quality estimates are insufficient. Interestingly, despite perfect post-hoc quality estimation for SWE-Bench, the baseline cascade strategy also performs poorly. This is due to the binary feedback of the quality estimator, which leads the threshold $\tau$ of the baseline cascade to either admit all models ($\tau=0$) or only correct ones ($\tau>0$).
For Minerva Math and LiveCodeBench, the opposite trend holds true. With accurate ex-ante quality estimation, the routing strategy achieves strong performance, surpassing the baseline cascade strategy by $10\%$. However, the cascade routing strategy still outperforms all methods, highlighting its robustness across diverse benchmarks and quality estimation scenarios. Interestingly, despite poor post-hoc quality estimation, our cascading strategy nearly matches the performance of routing. This suggests that the cascade effectively leverages ex-ante quality estimation to make informed decisions, unlike the baseline cascade.
Poor Quality Estimation
We perform experiments on classification and open-form reasoning tasks where there is no known accurate quality estimator. The classification benchmarks include ARC-Challenge (Clark et al., 2018), MMLU-Pro (Wang et al., 2024), and MixEval (Ni et al., 2024). For open-form reasoning tasks, we use MMLU-Pro and GSM8k (Cobbe et al., 2021). In classification, models select a single option representing their answer, with no intermediate reasoning process. In contrast, open-form reasoning allows models to generate their answers after reasoning. In this section, we evaluate two model families consisting of three models, Llama and Gemma, and show similar numbers for the Mistral model family in App. H. We create a quality estimator based on state-of-the-art work Gupta et al. (2024), which uses log probabilities as features. For full details on the benchmarks, models, and cost and quality estimators, we refer to App. D.
@@
Table 2 (right) presents the results for the Llama and Gemma model families across both benchmarks. Cascade routing consistently performs on par with or outperforms all baselines, though with much narrower margins reaching up to $1.2\%$. This reduced gain can be attributed to the fact that the quality and cost estimates are very noisy, leading to performance gains over the naive baseline similar to those observed in very high-noise scenarios on RouterBench.
@@ 6 Related Work
Routing is a widely studied problem in machine learning, particularly in the task of directing input queries to specialized models. One of the most common applications of routing is model selection for natural language input queries with a known answer (Chuang et al., 2024; Ding et al., 2024; Hari & Thomson, 2023; Liu et al., 2024; Jang et al., 2023; Nguyen et al., 2024; Sakota et al., 2024; Shnitzer et al., 2023). All these works train a model to predict whether a given model will correctly answer a query. Though the setups in these works are largely similar, they vary in certain specifics, such as the type of input queries or the features used for quality estimation.
Routing is also applied in other areas. For instance, Lu et al. (2024); Ong et al. (2024) use preference data to train a quality estimator, which facilitates routing in scenarios involving real-world user queries where clear ground-truth answers may not exist. Additionally, Chen et al. (2022) employ routing for API selection in multi-label classification tasks, focusing on directing queries to the appropriate API based on task requirements. Similarly, Zhang et al. (2024b) apply routing in software agent environments, directing user issues to the agent most suited to handle them. Finally, Pichlmeier et al. (2024) dynamically routes token generation instead of entire queries, allowing for more fine-grained routing decisions.
@@
Cascading techniques are primarily used to reduce inference costs by employing smaller models initially and only cascading to larger models if the smaller ones fail to provide a sufficiently accurate answer. Most often, cascading decisions are based on the smaller model’s confidence in its own predictions (Chen et al., 2023, 2024; Ramírez et al., 2024; Varshney & Baral, 2022). However, alternative techniques also exist. For example, Madaan et al. (2023) propose running models multiple times and measuring the variance in their responses to decide whether to cascade to a larger model.
@@
There has also been specific research on quality estimation within cascading frameworks. Gupta et al. (2024) examine various measures of uncertainty in language model answers, evaluating their impact on cascading performance. Meanwhile, Jitkrittum et al. (2023) explore failure cases in cascading mechanisms that rely on uncertainty, introducing alternative quality measures that enhance cascade efficiency. Furthermore, Xue et al. (2023) apply cascading to majority voting for a single model to obtain a method called dynamic voting: the cascade stops depending on the aggregated answers of all previous model computations. Lastly, (Zhang et al., 2024a) propose the use of multi-objective quality metrics to guide cascading decisions and do not solely focus on accuracy.